OpenStudy (anonymous):

How to simplify -5√18x Do you simplify the 5 since nothing will go into it? I know you can simplify √18x by doing something like: √9x * 2x = √9x * √2x = ?? I'm very confused!

jimthompson5910 (jim_thompson5910):

is it \[\Large -5\sqrt{18x}\] OR is it \[\Large -\sqrt[5]{18x}\]

OpenStudy (anonymous):

it's the first one

jimthompson5910 (jim_thompson5910):

ok you can break up 18x into 9*2x (kinda like what you're doing above) to get \[\Large -5\sqrt{18x}\] \[\Large -5\sqrt{9*2x}\] \[\Large -5\sqrt{9}*\sqrt{2x}\] \[\Large -5*3*\sqrt{2x}\] \[\Large -15\sqrt{2x}\]

jimthompson5910 (jim_thompson5910):

the only difference is that there isn't an x with the 9 since you can't factor x into x times x

OpenStudy (anonymous):

Oh okay, I see what you're saying, thank you!

jimthompson5910 (jim_thompson5910):

yw